2.7 moles of Tang are added to a pitcher containing 3.28 liters of H2O at a temperature of 25°C. The vapor pressure of water alone is 23.8 mm Hg at 25o C. What is the new vapor pressure of the Tang
solution?
1
Expert's answer
2020-04-08T13:25:57-0400

According to Rault's law, vapor pressure of water in a solution (p) is related to vapor pressure of pure water (p0) and molar fraction of the dissolved substance x (Tang is not electrolyte, so isotonic coefficient is unity) through the formular:

(p0 - p)/p0 = x

In the solution there are 2.7 moles of Tang and 3.28 (liter) *1000 (g/liter) / 18 (g/mole) = 182.2 moles of water. So, molar fraction of Tang is: x = 2.7/(2.7+182.2) = 0.014

p = p0*(1-x) = 23.8*(1-0.014) = 23.4668 = around 23.5 mm Hg